4. (a) Find and write down the general solution of the ODE 2y'-x^3=0 in the form of a power series about x = 0. Only include the first three non-zero terms in each of the two linearly independent solutions (in an interval I centered at x = 0) that you obtain. (b) Check that each of the two linearly independent solutions you found in part (a) individually satisfies the ODE, up through terms of order x^12
Answer #1

Now equation is Given that 0 24- x3y=0 re write the 24 - x3y=0 Let 9. E anzn be its pocoer Series solution then 2y- x3y=0a5 a6 = al 10 ao X Then ao ao 2.4.5 40 au 2.5.6-c0 Az az -0 (02-0) 2.6.7 ag=0 , ag=0 a5 a5 2.9.10 180 7200 AG A6 al 2.10.11NOW yait 75 y x16 ) 40 127200 y u6 Zoll 겐 0 + + 2 Go 13200 => y,- x4 u9 of 8 ym klo = 1+25 720 10 + 1200 yl x3 žtug yll 80 t